OpenStudy (anonymous):

Multiply: 3xsqrt(2) * sqrt(14x)? 6xsqrt(7x) 7xsqrt(6x) 12xsqrt(x) 21xsqrt(2x)

OpenStudy (anonymous):

\[\sqrt{A \times B} = \sqrt{A} \times \sqrt{B}\]

OpenStudy (anonymous):

6xsqrt(2x)

OpenStudy (anonymous):

that isnt one of my options

jimthompson5910 (jim_thompson5910):

It should be \[\large 6x\sqrt{7x}\]

jimthompson5910 (jim_thompson5910):

this is because the 14x factors into 2*7x

OpenStudy (anonymous):

oh okay thanks i tried this question alot and just couldnt get it